OpenStudy (anonymous):

Solve for x: -5(x - 3) = 30 x = -9 x = 3 x = 9 x = -3

ganeshie8 (ganeshie8):

start with dividing by "-5" both sides

ganeshie8 (ganeshie8):

-5(x - 3) = 30 x-3 = -6 x = ?

OpenStudy (anonymous):

\[-5\left( x-3 \right)=30\] \[-5x-15=30\] \[-5x+15=30\] +15 cross both 15 out and you should end up with -5x=45 divide both sides by -5 and 45/-5= x=-9 Is this Clear ?

OpenStudy (anonymous):

Yes, thank you guys :)

OpenStudy (anonymous):

welcome